# Representing Matroids over the Reals is $\exists\mathbb{R}$-complete

### v3

@@ 1 Introduction
Many articles on matroids assume that the matroid is representable [41, 16, 18]. Representability either heavily simplifies proofs and definitions or is even essential. We show that the question of representability over the reals is as difficult as the existential theory of the reals, that is $\exists\mathbb{R}$-complete. The complexity class $\exists\mathbb{R}$ can be defined as the family of algorithmic problems that is polynomial-time equivalent to determining if a multivariate polynomial (with integer coefficients) has a real root, see Section 2 for an introduction and overview of this complexity class.
@@
The rank ${\sf rk}(M)$ of a matroid $M$ is the maximum size of an independent set. We say that a matroid $M=(E,\mathcal{I})$ is representable over $\mathbb{F}$ if there is a matrix $A$ over $\mathbb{F}$ such that $M=M[A]$. Note that all columns of $A$ live in a subspace of dimension at most ${\sf rk}(M[A])$. Therefore, we can assume without loss of generality that the columns of $A$ have dimension ${\sf rk}(M[A])$. We refer to [51] for more background on matroids.
@@
An immediate question that arises from this picture is whether a picture exists where the circle is not used and the dependencies are all pictured by lines. This is equivalent to asking whether the Fano matroid is representable over the reals. It is well-known not to be [51, Proposition 6.4.8]. The problem Matroid $\mathbb{R}$-Representability addresses the general question of deciding whether a given matroid can be represented like that, and our main result is that this problem is $\exists\mathbb{R}$-complete.
@@
Theorem 1 easily follows from Theorem 2 as deciding whether an order type is representable over the reals is $\exists\mathbb{R}$-complete. This follows from techniques dating back to the proof of the Mnëv Universality Theorem (see [45, 53]). However, as explained in [45], the proof that stretchability is $\exists\mathbb{R}$-hard requires a significant number of intricate steps, some of which can be simplified in the setting of Matroid $\mathbb{R}$-Representability. Indeed, the need for different scales (see for example [45, Proof of Theorem 4.6]), which is the main difficulty in the oriented case, can be completely circumvented in our case. Therefore, for the sake of completeness and simplicity, we also provide a self-contained proof of Theorem 1. We think that it might be educational to first understand the proof of Theorem 1, before one tries to understand the $\exists\mathbb{R}$-completeness of stretchability.
@@ 2 Proof Overview and Background
While we could not find any prior proof of Theorem 1 in the literature (hence this work), there are many related works from at least two perspectives. First, as already mentioned, when one considers order-types instead of unoriented matroids, Theorem 1 is very well-known. Second, topological universality theorems have been proved for the real-representability of matroids in the algebraic geometry literature, see for example Lafforgue [39] and Lee and Vakil [40]. While our work uses tools that are similar in spirit to those papers, it differs in that our constructions are arguably simpler and that we specifically focus on proving the computational hardness result, which is not the point of focus of those previous works, and is not entirely equivalent (see discussion below). Furthermore, we believe that it is worthwhile to have a complete proof of Theorem 1 in a purely combinatorial language, as opposed to previous works which employ the language of schemes.
Shor’s proof of Mnëv’s universality theorem [60, Section 4] introduces an intermediate problem called the existential theory of totally real ordered variables, which is very similar to Distinct-ETR but features totally ordered variables $x_{1}<x_{2}\ldots<x_{n}$ as opposed to just requiring distinctness as in our problem. This ordering is desirable when one investigates oriented matroids, and unneeded for unoriented ones. This difference allows for a proof that is arguably simpler than his, or at least different.
@@
A practical reason why matroids are relevant for computational purposes is that they capture in a simple way the class of discrete objects where greedy algorithms are successful in finding an optimal solution. For example, the standard Kruskal and Prim algorithms to compute a Minimum Spanning Tree in a weighted graph can be abstracted by considering the vector matroid defined by an oriented incidence matrix of the graph (called a graphic matroid), and then generalized to compute in polynomial time a maximum or minimum-weight basis for any matroid. This property actually characterizes matroids, see for example Oxley [51, Section 1.8]
@@
For an algorithm on matroids, a suitable encoding scheme of the input matroid is needed. A common way is to take the input matroid $M=(E,\mathcal{I})$ in the form of an independence oracle which answers whether a given subset of the ground elements $E$ is independent or not. There are algorithms which run with a polynomial number of queries to such an oracle. For example, a maximum weight independent set of a given matroid can be computed in this way. However, for many natural matroid properties, it is known that there is no algorithm with polynomially bounded queries to an independence oracle [31] including representability over the finite field with two elements $GF(2)$ and connectivity of a matroid.
A vector representation of a matroid offers a compelling alternative to an independence oracle as matroid operations can be substantially more efficient using matrix operations. The Matroid parity problem, a common generalization of graph matching and matroid intersection, is solvable in polynomial time given a vector representation [41] while a super-polynomial number of calls is needed under the independence oracle model [31]. Deciding whether the branch-width of a matroid is at most $k$ is a common generalization of computing the branch-width, rank-width and carving-width of a graph. While there is an algorithm with $n^{O(k)}$ queries on an $n$ element matroid for this problem [50] under the oracle model, it remains unknown whether the dependency on $k$ in the exponent can be replaced by a uniform constant. In contrast, the branch-width of a vector matroid can be computed in $f(k)\cdot n^{3}$ time [32] when the given representation is over a finite field $\mathbb{F}$.
Another powerful application of a vector representation can be found in the theory of kernelization in parameterized complexity. A surprising discovery of Kratsch and Wahlströhm [38] is that for many graph cut problems, compressing the input boils down to finding a so-called representative set of a matroid. When the said matroid is a vector matroid, a representative set of bounded size can be efficiently computed in polynomial time [43, 44]. It turns out that solutions to graph cut problems can be encoded as independent sets in gammoids, which form a well-known class of representable matroids and of which a vector representation can be constructed in randomized polynomial time.
@@
One might wonder why we jumped from realizability of matroids to realizability of abstract order types, instead of using the perhaps closer notion of oriented matroids [12], for which one can also define a realizability problem and investigate its complexity. The reason is that in our arguments, we reason extensively with point configurations, and the geometric interpretation described above does not adapt directly to oriented matroids, as scaling a column by a negative number could lead to a change of the underlying oriented matroid. The correct framework to connect oriented matroids to point configurations is to only consider acyclic oriented matroids [12, Section 1.2.b], that is, those for which the geometric interpretation works readily without a need for rescaling by a negative number. This notion of acyclic, oriented matroids coincides with the notion of abstract order types, and so do their realizability problems. Note that in some of the existing literature, oriented matroids and abstract order types are sometimes described as equivalent. Therefore, we stress this subtle difference here.
@@
The complexity class $\exists\mathbb{R}$ (pronounced as ‘ER’, ‘exists R’, or ‘ETR’) has gained a lot of interest in recent years. It is defined via its canonical complete problem ETR (short for Existential Theory of the Reals. ETR refers to a geometric problem and $\exists\mathbb{R}$ refers to the complexity class. While there are several different variants of ETR, there is only one complexity class.) and contains all problems that polynomial-time many-to-one reduce to it. In an ETR instance, we are given a sentence of the form
@@
Here the first inclusion follows because a SAT instance can trivially be written as an equivalent ETR instance. The second inclusion is highly non-trivial and was first proven by Canny in his seminal paper [19].
Note that the complexity of working with continuous numbers was studied in various contexts. To avoid confusion, let us make some remarks on the underlying machine model. The underlying machine model for $\exists\mathbb{R}$ (over which sentences need to be decided and where reductions are performed) is the word RAM (or equivalently, a Turing machine) and not the real RAM [25] or the Blum-Shub-Smale model [13].
The complexity class $\exists\mathbb{R}$ gains its importance by numerous important algorithmic problems that have been shown to be complete for this class in recent years. The name $\exists\mathbb{R}$ was introduced by Schaefer in [53] who also pointed out that several NP-hardness reductions from the literature actually implied $\exists\mathbb{R}$-hardness. For this reason, several important $\exists\mathbb{R}$-completeness results had been obtained before the need for a dedicated complexity class became apparent.
Common features of $\exists\mathbb{R}$-complete problems are their continuous solution space and the nonlinear relations between their variables. Important $\exists\mathbb{R}$-completeness results include the realizability of abstract order types [49, 60] and geometric linkages [54], as well as the recognition of geometric segment [37, 45], unit-disk [35, 47], and ray intersection graphs [20]. More results appeared in the graph drawing community [23, 24, 42, 55], regarding the Hausdorff distance [34], regarding polytopes [22, 52], the study of Nash-equilibria [7, 10, 11, 26, 56], training neural networks [3, 9], matrix factorization [21, 57, 58, 59, 62], or continuous constraint satisfaction problems [48]. In computational geometry, we would like to mention geometric packing [5], the art gallery problem [2], and covering polygons with convex polygons [1].
Recall that NP is usually described using a witness and a verification algorithm. The same characterization exists for $\exists\mathbb{R}$. Instead of the witness consisting of binary words of polynomial length, we allow in addition using real-valued numbers as a witness. Furthermore, in order to be able to use those real numbers, we are allowed to work on the so-called real RAM model of computation. The real RAM allows arithmetic operations with real numbers in constant time [25].
@@
Many results and techniques on the existential theory of the reals actually precede the study of this complexity class. The underlying idea was to study how complicated solution spaces can be from a topological perspective. For example, if we want to study convex polytopes, we are often interested in the properties of their face lattice, which is the family of faces of different dimension together with their inclusion order The face lattice is a purely combinatorial object. Therefore, it is natural to ask which face lattices are realizable by polytopes. If there existed an easy combinatorial description of realizable face lattices, convex polytopes would be much better understood. Given a specific face lattice $L$, we can study its suitably defined solution space $S(L)$. As the realizability question can be formulated as an ETR-formula, it follows that $S(L)$ is a semi-algebraic set. Now, let $T$ be a different semi-algebraic set, we wonder whether there exists a face lattice $L$ such that $S(L)$ is homotopy-equivalent to $T$. Maybe surprisingly topological universality states that there is such an $L$ for any semi-algebraic set $T$. This type of property feels very strong, as it intuitively states that we can encode the vast complexity of semi-algebraic sets into the problem of realizing convex polytopes. Indeed, many of the results that establish such topological universality also imply $\exists\mathbb{R}$-completeness [45].
However, topological universality can also be established for NP-complete problems as has been shown by Bertschinger, El Maalouly, Miltzow, Schnider, and Weber [8]. As they showed, for simplicial complexes, it is sufficient to encode the topology as it is possible to triangulate semi-algebraic sets [29]. In other words, the difference between the wild semi-algebraic sets and the tame simplicial complexes is not that the former emit a more complicated topological structure. The difference comes from the ability of semi-algebraic sets to encode complicated topological spaces in a much more concise manner. (To be precise the description complexity of a topological space might be exponentially smaller using the language of semi-algebraic sets, compared to simplicial complexes.) $\exists\mathbb{R}$-completeness can be interpreted as giving a concise encoding of semi-algebraic sets into a different domain. To make our life easier, we do not care about preserving the complete topology, but merely the property of being empty or not. Still, in order to do so one usually also preserves topological properties. This is the reason why there is a close connection between topological universality and $\exists\mathbb{R}$-completeness. But given that NP-complete problems may also admit universality theorems $\exists\mathbb{R}$-completeness may arguably be considered the more interesting finding.
@@
We want to point out that previous universality results often also showed stronger results than mere topological universality. For instance, Richter-Gebert showed such a stronger universality theorem for polytopes [52]. Given a face lattice $F$, we can define the set of polytopes $V(F)$ having face lattice $F$. Richter-Gebert showed that for every semi-algebraic set $S$ there exists a face-lattice $F$ of a polytope such that $V(F)$ is stably-equivalent to $S$. To define the notion of stable-equivalence goes above the scope of this paper. We just note that stable-equivalence encapsulates more than just the topology of $S$, but also to a degree the “geometry” of $S$.
@@
Recall that an input graph for a graph problem can be given as an adjacency matrix, or equivalently the family of vertex subsets of size two. Similarly, an input hypergraph is typically given as a set family with an explicit description of all hyperedges. An immediate analogue of such an hypergraph description for a matroid is an explicit enumeration of all bases (maximal independent sets) or all circuits (minimally dependent sets). However, explicitly stating all independent sets, bases, or circuits is unconventional; the most common size measure of a matroid is the number of elements, which is polynomially bounded by the size of a graph or matrix that is generalized by a matroid. On the other hand, the number of bases or circuits can be prohibitively large in comparison to the number of elements. Specifically, it is known that the number of distinct matroids on $n$ elements is doubly exponential in $n$ [36], hence in space of size polynomial in $n$ one cannot describe an arbitrary input matroid. For further details about explicit matroid encoding, see [46].
@@
Due to the limitations of the above two explicit descriptions, the most common way to encode an input matroid without any restriction is with an oracle, often an independence oracle. One can view an independence oracle as a black box expressing a boolean function on $n$ variables. The boolean function is on $n$ input variables and outputs 0 or 1 depending on whether the input corresponds to (a characteristic vector of) an independent set of the said matroid. Problems with black box functions, e.g., an implicit input with oracle access are studied in the context of learning a function with a small number of queries, e.g. Polynomial Identity Testing, and also in the context of search problems where a graph is accessed by adjacency query. Such a problem does not fit in the classic computational complexity, where an explicit string of numbers is expected as an input. Moreover, learning a black box (boolean) function cannot be done efficiently. As mentioned previously, even when the boolean input functions are restricted to be matroid oracles, deciding whether a nontrivial matroid property holds or not requires $2^{\Theta(n)}$ queries [61] even for basic properties such as connectivity and representability over $\mathbb{F}=GF(2)$, and even with a randomized algorithm.
@@
The presented verification algorithm shows that $\mathbb{F}$-representability is in NP for each finite field $\mathbb{F}.$ A matrix over $\mathbb{F}$ as witness and the polynomial-time verification algorithm naturally extend to the case when $\mathbb{F}=\mathbb{R}$, where the verification algorithm works on a real RAM. Therefore, $\mathbb{R}$-representability is in $\exists\mathbb{R}$. For details on the characterization of $\exists\mathbb{R}$ via a witness and a polynomial-time verification algorithm on a real RAM, see [25], and also the discussion in the paragraph above about the existential theory of the reals.
@@
Furthermore, it is likely that $\mathbb{F}$-representability is in co-NP for each finite field $\mathbb{F}.$ It is known [28] that for any prime field $\mathbb{F}$, non-$\mathbb{F}$-representability can be certified by evaluating the ranks of $O(n^{2})$ subsets of an $n$-element matroid. Notice that when the matroid $M=(E,\mathcal{B})$ is given with an explicit description of all the bases $\mathcal{B}$ of $M$, evaluating ${\sf rk}(X)$ for $X\subseteq E$ can be done in time polynomial in the input size because ${\sf rk}(X)$ equals the maximum of $|X\cap B|$ over all $B\in\mathcal{B}.$ Therefore, $\mathbb{F}$-representability is in co-NP under the explicit bases description for each prime field $\mathbb{F}.$ For an arbitrary finite field $\mathbb{F}$, not necessarily prime, up to our best knowledge there is no published result which establishes that a polynomial number of rank evaluations suffices for non-$\mathbb{F}$-representability. However, it is known that a positive resolution of Rota’s conjecture implies that only a constant, depending on $|\mathbb{F}|$ only, number of rank evaluations would suffice  [51] to certify that a given matroid is not $\mathbb{F}$-representable. The proof of Rota’s conjecture was announced in 2014 by Geelen, Gerards and Whittle [27] although it is expected to take a few more years for the full proof to be written for publication.

### v4

@@ Abstract
keywords:
Computer Science - Computational Complexity, Mathematics - Combinatorics
@@ 1 Introduction
Many articles on matroids assume that the matroid is representable, see for example Lovász (1980); Cameron, Dinu, Michałek, and Seynnaeve (2022); Cameron and Mayhew (2021). Representability either heavily simplifies proofs and definitions or is even essential. We show that the question of representability over the reals is as difficult as the existential theory of the reals, that is $\exists\mathbb{R}$-complete. The complexity class $\exists\mathbb{R}$ can be defined as the family of algorithmic problems that is polynomial-time equivalent to determining if a multivariate polynomial (with integer coefficients) has a real root, see Section 2 for an introduction and overview of this complexity class.
@@
The rank ${\sf rk}(M)$ of a matroid $M$ is the maximum size of an independent set. We say that a matroid $M=(E,\mathcal{I})$ is representable over $\mathbb{F}$ if there is a matrix $A$ over $\mathbb{F}$ such that $M=M[A]$. Note that all columns of $A$ live in a subspace of dimension at most ${\sf rk}(M[A])$. Therefore, we can assume without loss of generality that the columns of $A$ have dimension ${\sf rk}(M[A])$. We refer to Oxley (2006) for more background on matroids.
@@
An immediate question that arises from this picture is whether a picture exists where the circle is not used and the dependencies are all pictured by lines. This is equivalent to asking whether the Fano matroid is representable over the reals. It is well-known not to be (Oxley, 2006, Proposition 6.4.8). The problem Matroid $\mathbb{R}$-Representability addresses the general question of deciding whether a given matroid can be represented like that, and our main result is that this problem is $\exists\mathbb{R}$-complete.
@@
Theorem 1 easily follows from Theorem 2 as deciding whether an order type is representable over the reals is $\exists\mathbb{R}$-complete. This follows from techniques dating back to the proof of the Mnëv Universality Theorem (Matoušek, 2014; Schaefer, 2010). However, as explained in Matoušek (2014), the proof that stretchability is $\exists\mathbb{R}$-hard requires a significant number of intricate steps, some of which can be simplified in the setting of Matroid $\mathbb{R}$-Representability. Indeed, the need for different scales (see for example (Matoušek, 2014, Proof of Theorem 4.6)), which is the main difficulty in the oriented case, can be completely circumvented in our case. Therefore, for the sake of completeness and simplicity, we also provide a self-contained proof of Theorem 1. We think that it might be educational to first understand the proof of Theorem 1, before one tries to understand the $\exists\mathbb{R}$-completeness of stretchability.
@@ 2 Proof Overview and Background
While we could not find any prior proof of Theorem 1 in the literature (hence this work), there are many related works from at least two perspectives. First, as already mentioned, when one considers order-types instead of unoriented matroids, Theorem 1 is very well-known. Second, topological universality theorems have been proved for the real-representability of matroids in the algebraic geometry literature, see for example Lafforgue (2003) and Lee and Vakil (2013). While our work uses tools that are similar in spirit to those papers, it differs in that our constructions are arguably simpler and that we specifically focus on proving the computational hardness result, which is not the point of focus of those previous works, and is not entirely equivalent (see discussion below). Furthermore, we believe that it is worthwhile to have a complete proof of Theorem 1 in a purely combinatorial language, as opposed to previous works which employ the language of schemes.
Shor’s proof of Mnëv’s universality theorem (Shor, 1991, Section 4) introduces an intermediate problem called the existential theory of totally real ordered variables, which is very similar to Distinct-ETR but features totally ordered variables $x_{1}<x_{2}\ldots<x_{n}$ as opposed to just requiring distinctness as in our problem. This ordering is desirable when one investigates oriented matroids, and unneeded for unoriented ones. This difference allows for a proof that is arguably simpler than his, or at least different.
@@
A practical reason why matroids are relevant for computational purposes is that they capture in a simple way the class of discrete objects where greedy algorithms are successful in finding an optimal solution. For example, the standard Kruskal and Prim algorithms to compute a Minimum Spanning Tree in a weighted graph can be abstracted by considering the vector matroid defined by an oriented incidence matrix of the graph (called a graphic matroid), and then generalized to compute in polynomial time a maximum or minimum-weight basis for any matroid. This property actually characterizes matroids, see for example (Oxley, 2006, Section 1.8)
@@
For an algorithm on matroids, a suitable encoding scheme of the input matroid is needed. A common way is to take the input matroid $M=(E,\mathcal{I})$ in the form of an independence oracle which answers whether a given subset of the ground elements $E$ is independent or not. There are algorithms which run with a polynomial number of queries to such an oracle. For example, a maximum weight independent set of a given matroid can be computed in this way. However, for many natural matroid properties, it is known that there is no algorithm with polynomially bounded queries to an independence oracle (Jensen and Korte, 1982) including representability over the finite field with two elements $GF(2)$ and connectivity of a matroid.
A vector representation of a matroid offers a compelling alternative to an independence oracle as matroid operations can be substantially more efficient using matrix operations. The Matroid parity problem, a common generalization of graph matching and matroid intersection, is solvable in polynomial time given a vector representation (Lovász, 1980) while a super-polynomial number of calls is needed under the independence oracle model (Jensen and Korte, 1982). Deciding whether the branch-width of a matroid is at most $k$ is a common generalization of computing the branch-width, rank-width and carving-width of a graph. While there is an algorithm with $n^{O(k)}$ queries on an $n$ element matroid for this problem (Oum and Seymour, 2007) under the oracle model, it remains unknown whether the dependency on $k$ in the exponent can be replaced by a uniform constant. In contrast, the branch-width of a vector matroid can be computed in $f(k)\cdot n^{3}$ time (Jeong, Kim, and Oum, 2021) when the given representation is over a finite field $\mathbb{F}$.
Another powerful application of a vector representation can be found in the theory of kernelization in parameterized complexity. A surprising discovery of Kratsch and Wahlströhm (Kratsch and Wahlström, 2020) is that for many graph cut problems, compressing the input boils down to finding a so-called representative set of a matroid. When the said matroid is a vector matroid, a representative set of bounded size can be efficiently computed in polynomial time (László, 1977; Marx, 2009). It turns out that solutions to graph cut problems can be encoded as independent sets in gammoids, which form a well-known class of representable matroids and of which a vector representation can be constructed in randomized polynomial time.
@@
One might wonder why we jumped from realizability of matroids to realizability of abstract order types, instead of using the perhaps closer notion of oriented matroids (Björner, Las Vergnas, Sturmfels, White, and Ziegler, 1999), for which one can also define a realizability problem and investigate its complexity. The reason is that in our arguments, we reason extensively with point configurations, and the geometric interpretation described above does not adapt directly to oriented matroids, as scaling a column by a negative number could lead to a change of the underlying oriented matroid. The correct framework to connect oriented matroids to point configurations is to only consider acyclic oriented matroids (Björner, Las Vergnas, Sturmfels, White, and Ziegler, 1999, Section 1.2.b), that is, those for which the geometric interpretation works readily without a need for rescaling by a negative number. This notion of acyclic, oriented matroids coincides with the notion of abstract order types, and so do their realizability problems. Note that in some of the existing literature, oriented matroids and abstract order types are sometimes described as equivalent. Therefore, we stress this subtle difference here.
@@
The complexity class $\exists\mathbb{R}$ (pronounced as ‘ER’, ‘exists R’, or ‘ETR’) has gained a lot of interest in recent years. See the compendium by Cardinal, Miltzow, and Schaefer (2024) for a comprehensive overview. The complexity class is defined via its canonical complete problem ETR (short for Existential Theory of the Reals. ETR refers to a geometric problem and $\exists\mathbb{R}$ refers to the complexity class. While there are several different variants of ETR, there is only one complexity class.) and contains all problems that polynomial-time many-to-one reduce to it. In an ETR instance, we are given a sentence of the form
@@
Here the first inclusion follows because a SAT instance can trivially be written as an equivalent ETR instance. The second inclusion is highly non-trivial and was first proven by Canny in his seminal paper (Canny, 1988).
Note that the complexity of working with continuous numbers was studied in various contexts. To avoid confusion, let us make some remarks on the underlying machine model. The underlying machine model for $\exists\mathbb{R}$ (over which sentences need to be decided and where reductions are performed) is the word RAM (or equivalently, a Turing machine) and not the real RAM (Erickson, van der Hoog, and Miltzow, 2020) or the Blum-Shub-Smale model (Blum, Shub, and Smale, 1989).
The complexity class $\exists\mathbb{R}$ gains its importance by numerous important algorithmic problems that have been shown to be complete for this class in recent years. The name $\exists\mathbb{R}$ was introduced in Schaefer (2010) who also pointed out that several NP-hardness reductions from the literature actually implied $\exists\mathbb{R}$-hardness. For this reason, several important $\exists\mathbb{R}$-completeness results had been obtained before the need for a dedicated complexity class became apparent.
Common features of $\exists\mathbb{R}$-complete problems are their continuous solution space and the nonlinear relations between their variables. Important $\exists\mathbb{R}$-completeness results include the realizability of abstract order types (Mnëv, 1988; Shor, 1991) and geometric linkages (Schaefer, 2013), as well as the recognition of geometric segment (Kratochvíl and Matoušek, 1994; Matoušek, 2014), unit-disk (Kang and Müller, 2012; McDiarmid and Müller, 2013), and ray intersection graphs (Cardinal, Felsner, Miltzow, Tompkins, and Vogtenhuber, 2018). More results appeared in the graph drawing community (Dobbins, Kleist, Miltzow, and Rzążewski, 2023; Erickson, 2019; Lubiw, Miltzow, and Mondal, 2022; Schaefer, 2021), regarding the Hausdorff distance (Jungeblut, Kleist, and Miltzow, 2022), regarding polytopes (Dobbins, Holmsen, and Miltzow, 2019; Richter-Gebert and Ziegler, 1995), the study of Nash-equilibria (Berthelsen and Hansen, 2019; Bilò and Mavronicolas, 2016, 2017; Garg, Mehta, Vazirani, and Yazdanbod, 2018; Schaefer and Štefankovič, 2017), training neural networks (Abrahamsen, Kleist, and Miltzow, 2021; Bertschinger, Hertrich, Jungeblut, Miltzow, and Weber, 2022), matrix factorization (Chistikov, Kiefer, Marusic, Shirmohammadi, and Worrell, 2016; Schaefer and Štefankovič, 2018; Shitov, 2016, 2017; Tuncel, Vavasis, and Xu, 2022), or continuous constraint satisfaction problems (Miltzow and Schmiermann, 2022). In computational geometry, we would like to mention geometric packing (Abrahamsen, Miltzow, and Seiferth, 2020), the art gallery problem (Abrahamsen, Adamaszek, and Miltzow, 2018), and covering polygons with convex polygons (Abrahamsen, 2022).
Recall that NP is usually described using a witness and a verification algorithm. The same characterization exists for $\exists\mathbb{R}$. Instead of the witness consisting of binary words of polynomial length, we allow in addition using real-valued numbers as a witness. Furthermore, in order to be able to use those real numbers, we are allowed to work on the so-called real RAM model of computation. The real RAM allows arithmetic operations with real numbers in constant time (Erickson, van der Hoog, and Miltzow, 2020).
@@
Many results and techniques on the existential theory of the reals actually precede the study of this complexity class. The underlying idea was to study how complicated solution spaces can be from a topological perspective. For example, if we want to study convex polytopes, we are often interested in the properties of their face lattice, which is the family of faces of different dimension together with their inclusion order The face lattice is a purely combinatorial object. Therefore, it is natural to ask which face lattices are realizable by polytopes. If there existed an easy combinatorial description of realizable face lattices, convex polytopes would be much better understood. Given a specific face lattice $L$, we can study its suitably defined solution space $S(L)$. As the realizability question can be formulated as an ETR-formula, it follows that $S(L)$ is a semi-algebraic set. Now, let $T$ be a different semi-algebraic set, we wonder whether there exists a face lattice $L$ such that $S(L)$ is homotopy-equivalent to $T$. Maybe surprisingly topological universality states that there is such an $L$ for any semi-algebraic set $T$. This type of property feels very strong, as it intuitively states that we can encode the vast complexity of semi-algebraic sets into the problem of realizing convex polytopes. Indeed, many of the results that establish such topological universality also imply $\exists\mathbb{R}$-completeness (Matoušek, 2014).
However, topological universality can also be established for NP-complete problems as has been shown by Bertschinger, El Maalouly, Miltzow, Schnider, and Weber (2023). As they showed, for simplicial complexes, it is sufficient to encode the topology as it is possible to triangulate semi-algebraic sets (Hironaka, 1975). In other words, the difference between the wild semi-algebraic sets and the tame simplicial complexes is not that the former emit a more complicated topological structure. The difference comes from the ability of semi-algebraic sets to encode complicated topological spaces in a much more concise manner. (To be precise the description complexity of a topological space might be exponentially smaller using the language of semi-algebraic sets, compared to simplicial complexes.) $\exists\mathbb{R}$-completeness can be interpreted as giving a concise encoding of semi-algebraic sets into a different domain. To make our life easier, we do not care about preserving the complete topology, but merely the property of being empty or not. Still, in order to do so one usually also preserves topological properties. This is the reason why there is a close connection between topological universality and $\exists\mathbb{R}$-completeness. But given that NP-complete problems may also admit universality theorems $\exists\mathbb{R}$-completeness may arguably be considered the more interesting finding.
@@
We want to point out that previous universality results often showed stronger results than mere topological universality. For instance, Richter-Gebert and Ziegler (1995) showed such a stronger universality theorem for polytopes. Given a face lattice $F$, we can define the set of polytopes $V(F)$ having face lattice $F$. Richter-Gebert showed that for every semi-algebraic set $S$ there exists a face-lattice $F$ of a polytope such that $V(F)$ is stably-equivalent to $S$. To define the notion of stable-equivalence goes above the scope of this paper. We just note that stable-equivalence encapsulates more than just the topology of $S$, but also to a degree the “geometry” of $S$.
@@
Recall that an input graph for a graph problem can be given as an adjacency matrix, or equivalently the family of vertex subsets of size two. Similarly, an input hypergraph is typically given as a set family with an explicit description of all hyperedges. An immediate analogue of such an hypergraph description for a matroid is an explicit enumeration of all bases (maximal independent sets) or all circuits (minimally dependent sets). However, explicitly stating all independent sets, bases, or circuits is unconventional; the most common size measure of a matroid is the number of elements, which is polynomially bounded by the size of a graph or matrix that is generalized by a matroid. On the other hand, the number of bases or circuits can be prohibitively large in comparison to the number of elements. Specifically, it is known that the number of distinct matroids on $n$ elements is doubly exponential in $n$ (Knuth, 1974), hence in space of size polynomial in $n$ one cannot describe an arbitrary input matroid. For further details about explicit matroid encoding, see Mayhew (2008).
@@
Due to the limitations of the above two explicit descriptions, the most common way to encode an input matroid without any restriction is with an oracle, often an independence oracle. One can view an independence oracle as a black box expressing a boolean function on $n$ variables. The boolean function is on $n$ input variables and outputs 0 or 1 depending on whether the input corresponds to (a characteristic vector of) an independent set of the said matroid. Problems with black box functions, e.g., an implicit input with oracle access are studied in the context of learning a function with a small number of queries, e.g. Polynomial Identity Testing, and also in the context of search problems where a graph is accessed by adjacency query. Such a problem does not fit in the classic computational complexity, where an explicit string of numbers is expected as an input. Moreover, learning a black box (boolean) function cannot be done efficiently. As mentioned previously, even when the boolean input functions are restricted to be matroid oracles, deciding whether a nontrivial matroid property holds or not requires $2^{\Theta(n)}$ queries (Truemper, 1982) even for basic properties such as connectivity and representability over $\mathbb{F}=GF(2)$, and even with a randomized algorithm.
@@
The presented verification algorithm shows that $\mathbb{F}$-representability is in NP for each finite field $\mathbb{F}.$ A matrix over $\mathbb{F}$ as witness and the polynomial-time verification algorithm naturally extend to the case when $\mathbb{F}=\mathbb{R}$, where the verification algorithm works on a real RAM. Therefore, $\mathbb{R}$-representability is in $\exists\mathbb{R}$. For details on the characterization of $\exists\mathbb{R}$ via a witness and a polynomial-time verification algorithm on a real RAM, see Erickson, van der Hoog, and Miltzow (2020), and also the discussion in the paragraph above about the existential theory of the reals.
@@
Furthermore, it is likely that $\mathbb{F}$-representability is in co-NP for each finite field $\mathbb{F}.$ It is known (Geelen and Whittle, 2016) that for any prime field $\mathbb{F}$, non-$\mathbb{F}$-representability can be certified by evaluating the ranks of $O(n^{2})$ subsets of an $n$-element matroid. Notice that when the matroid $M=(E,\mathcal{B})$ is given with an explicit description of all the bases $\mathcal{B}$ of $M$, evaluating ${\sf rk}(X)$ for $X\subseteq E$ can be done in time polynomial in the input size because ${\sf rk}(X)$ equals the maximum of $|X\cap B|$ over all $B\in\mathcal{B}.$ Therefore, $\mathbb{F}$-representability is in co-NP under the explicit bases description for each prime field $\mathbb{F}.$ For an arbitrary finite field $\mathbb{F}$, not necessarily prime, up to our best knowledge there is no published result which establishes that a polynomial number of rank evaluations suffices for non-$\mathbb{F}$-representability. However, it is known that a positive resolution of Rota’s conjecture implies that only a constant, depending on $|\mathbb{F}|$ only, number of rank evaluations would suffice (Oxley, 2006) to certify that a given matroid is not $\mathbb{F}$-representable. The proof of Rota’s conjecture was announced in 2014 by Geelen, Gerards, and Whittle (2014) although it is expected to take a few more years for the full proof to be written for publication.